OpenStudy (anonymous):

Which best describes the relationship between (x + 1) and the polynomial -3x3 - 2x2 + 1? A. (x + 1) is not a factor. B. It is impossible to tell if (x + 1) is a factor. C. (x + 1) is a factor. D. If (x + 1) isn't a factor, then x = -1 could be.

OpenStudy (anonymous):

by plugging x=-1 in -3x3 - 2x2 + 1 we have -3(-1)^3 -2(-1)^2 +1 =3-2+1=2<>0 hence A. (x + 1) is not a factor.